OpenStudy (anonymous):

HELP! I GIVE MEDALS:) rotate TRiangle ABC counterclockwise 270deg. A(2 -1) B(-2 1) C(1 3)

OpenStudy (anonymous):

@touseii45 do u think u could help me??

OpenStudy (anonymous):

is there a picture?

OpenStudy (anonymous):

I can draw one give me a sec

OpenStudy (anonymous):



OpenStudy (anonymous):

which point is it turning on?

OpenStudy (anonymous):

it doesn't say it just said find the matrix that results from a 120deg counterclockwise totation from triangle ABC

OpenStudy (anonymous):

okay uhm hold on(:

OpenStudy (anonymous):

kkk ill be here

OpenStudy (anonymous):

okay well first, do you know the definition of a what a matrix in this question is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

aha well your supposed to know that. i dont remember it

OpenStudy (anonymous):

ooh that part doesn't matter in the question lol i found what it means

OpenStudy (anonymous):

so do you know how to find the answer?

OpenStudy (anonymous):

ya i no who do find that i just need the coordinate of triangle ABC rotated 270deg counterclockwise

OpenStudy (anonymous):

maybe @Mertsj can help you lol

OpenStudy (anonymous):

or @jim_thompson5910

OpenStudy (anonymous):

@HawkCrimson can u help me out plezz :)

OpenStudy (anonymous):

@electrokid could u help me

OpenStudy (anonymous):

yup i got that

OpenStudy (anonymous):

oh you got it?

OpenStudy (anonymous):

what sems to be the problem?

OpenStudy (anonymous):

i got -1 as well

OpenStudy (anonymous):

no problem just cheeking my work with urs

OpenStudy (anonymous):

ok, so what was your rotation matrix?

OpenStudy (anonymous):

wait i think i got confused

OpenStudy (anonymous):

\[ A'=\left[\begin{matrix}0&1\\-1&0\end{matrix}\right]\left[\begin{matrix}2\\-1\end{matrix}\right]=\left[\begin{matrix}-1\\-2\end{matrix}\right] \]

OpenStudy (anonymous):

similarly for the other points

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is that what you got for A'?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

good

OpenStudy (anonymous):

so then is B 1 2

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

and C 3 -1

OpenStudy (anonymous):

bingo

OpenStudy (anonymous):

how about them medals now?

OpenStudy (anonymous):

sorry my sister had to use the computer for a sec thatnks for the help